[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 17, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 17 13:09:07 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
1. Graphicasts looked good tonight. Atlantic desk composed theirs on  
STARFISH and East Pacific desk's was completed on GRUNT.

2. Scatterometer Winds not available NAWIPS. Doug is working on this.
HSU:
No new issues.
Other items:
Marc is working with Verizon to fix the forwarding problem with the phones.  
All the extensions are working properly, but they are not being forwarded.  
Marc will update accordingly.

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171017/81931067/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list